OpenStudy (anonymous):

what is the distance from (6 -8) to the origin

OpenStudy (anonymous):

use the distance formula

OpenStudy (anonymous):

use pythogorean theorem. Or better yet the distance formula. origin = 0,0

OpenStudy (anonymous):

Distance to any point to another is given by \[D(AB)=\sqrt{(B_x - A_x)^2+(B_y - A_y)^2}\] A=(0,0) B=(6,-8) Paff.

OpenStudy (anonymous):



OpenStudy (kainui):

Is there a difference between the distance formula and the pythagorean theorem?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i got it thanks

OpenStudy (kainui):

Not really, it's actually the same idea.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

pythogorean theorem is used to derive distance formula